### v1

# Inversion Diameter and Treewidth

Abstract
In an oriented graph $\overrightarrow{G}$, the inversion of a subset $X$ of vertices consists in reversing the orientation of all arcs with both end-vertices in $X$. The inversion graph of a graph $G$, denoted by $\mathcal{I}(G)$, is the graph whose vertices are orientations of $G$ in which two orientations $\overrightarrow{G_{1}}$ and $\overrightarrow{G_{2}}$ are adjacent if and only if there is an inversion $X$ transforming $\overrightarrow{G_{1}}$ into $\overrightarrow{G_{2}}$. The inversion diameter of a graph $G$ is the diameter of its inversion graph $\mathcal{I}(G)$ denoted by $diam(\mathcal{I}(G))$. Havet, Hörsch, and Rambaud (2024) first proved that for $G$ of treewidth $k$, $diam(\mathcal{I}(G))\leq 2k$, and there are graphs of treewidth $k$ with inversion diameter $k+2$. In this paper, we construct graphs of treewidth $k$ with inversion diameter $2k$, which implies that the previous upper bound $diam(\mathcal{I}(G))\leq 2k$ is tight. Moreover, for graphs with maximum degree $\Delta$, Havet, Hörsch, and Rambaud (2024) proved $diam(\mathcal{I}(G))\leq 2\Delta-1$ and conjectured that $diam(\mathcal{I}(G))\leq\Delta$. We prove the conjecture when $\Delta=3$ with the help of computer calculations.
Keywords: inversion diameter; orientation; treewidth.
1 Introduction
An orientation of an undirected graph is an assignment of a direction to each edge, turning the initial graph into a directed graph. Let $G$ be a simple graph and $\overrightarrow{G_{1}}$ an orientation of $G$. If $X$ is a vertex set of $G$, the inversion of $X$ on $\overrightarrow{G_{1}}$ is a orientation $\overrightarrow{G_{2}}$ by reversing the orientation of all arcs with both ends in $X$.
The concept of inversion was first introduced by Belkhechine et al. [4]. They studied the inversion number of a directed graph $D$, denoted by $inv(D)$, which is the minimum number of inversions that transform $D$ into an acyclic graph. They proved, for every fixed $k$, given a tournament $T$, determining whether $inv(T)\leq k$ is polynomial-time solvable. In contrast, Bang-Jensen et al. [3] proved that given any directed graph $D$, determining whether $inv(D)\leq 1$ is NP-complete.
The maximum inversion numbers of all oriented graphs of order $n$, denoted by $inv(n)$, has also been investigated. Aubian et al. [2] and Alon et al. [1] proved $n-2\sqrt{2\log n}\leq inv(n)\leq n-\lceil\log(n+1)\rceil$. Besides these results, various related questions have also been studied.
Let $G$ be a simple graph. The inversion is a transformation between different orientations of $G$. Instead of transforming an orientation into an acyclic orientation, it is also natural to consider the transformation between two orientations. The inversion graph of $G$ denoted by $\mathcal{I}(G)$, is the graph whose vertices are the orientations of $G$ in which two orientations $\overrightarrow{G_{1}}$ and $\overrightarrow{G_{2}}$ are adjacent if and only if there is an inversion $X$ transforming $\overrightarrow{G_{1}}$ into $\overrightarrow{G_{2}}$. The inversion diameter of $G$ is the diameter of $\mathcal{I}(G)$, denoted by $diam(\mathcal{I}(G))$. It represents the maximum number of required inversions to transform an orientation of $G$ into another orientation of it.
Havet et al. [5] first introduced inversion diameter and studied it on various class of graphs. Let $G$ be a graph and let $<$ be a total ordering on $V(G)$. For every pair $u,u^{\prime}$ of vertices in $G$, let $N_{<u^{\prime}}(u)=\{v\in N(u)\mid v<u^{\prime}\}$ and $N_{>u^{\prime}}(u)=\{v\in N(u)\mid v>u^{\prime}\}$. We simply write $N_{<}(u)$ for $N_{<u}(u)$ and $N_{>}(u)$ for $N_{>u}(u)$. The ordering $<$ is $t$-strong if for every $u\in V(G)$
  • $|N_{<}(u)|+\log(|\{X\subseteq V(G)\mid\exists v\in N_{>}(u),X\subseteq N_{<u}(v)\}|)<t$, if $N_{>}(u)\neq\emptyset$, and
  • $N_{<}(u)\leq t$ otherwise.
A graph is strongly $t$-degenerate if it admits a $t$-strong ordering of its vertices. Havet et al. [5] showed that
Theorem 1.1 (Havet et al. [5])
Let $G$ be a graph and let $t$ be a positive integer. If $G$ is strong $t$-degenerate, then $diam(\mathcal{I}(G))\leq t$.
With the help of Theorem 1.1, they showed various bounds on $diam(\mathcal{I}(G))$ depending on the structure of $G$ as following:
Theorem 1.2 (Havet et al. [5])
  1. 1.
    For every graph $G$ with at least one edge and maximum degree $\Delta$, $diam(\mathcal{I}(G))\leq 2\Delta-1$.
  2. 2.
    $diam(\mathcal{I}(G))\leq 12$ for every planar graph $G$.
  3. 3.
    $diam(\mathcal{I}(G))\leq 2k$ for ever graph $G$ of treewidth at most $k$.
Havet et al. [5] also proved that for fixed $k\geq 2$, given a graph $G$, determining whether $diam(\mathcal{I}(G))\leq k$ is NP-hard. For a graph $G$ with maximum degree $3$ (sub-cubic graph), Havet et al. [5] showed a better bound $diam(\mathcal{I}(G))\leq 4$. Moreover, they proposed a conjecture on graphs with maximum degree $\Delta$ as Conjecture 1.3.
Conjecture 1.3 (Havet et al. [5])
For every graph $G$ with at least one edge and maximum degree $\Delta$, $diam(\mathcal{I}(G))\leq\Delta$.
The conjecture is true for $\Delta\leq 2$ [5]. In this paper, we prove the conjecture when $\Delta=3$. Computer assistance will be used in the proof of Theorem 1.4. A pure mathematical proof is still worth studying.
Theorem 1.4
If $G$ is a graph of maximum degree $3$, then $diam(\mathcal{I}(G))\leq 3$.
For graphs with treewidth at most $k$, Havet et al. [5] showed that there are graphs of treewidth at most $k$ with inversion diameter $k+2$. In this paper, we show that the upper bound $diam(\mathcal{I}(G))\leq 2k$ for graphs of treewidth at most $k$ is tight by proving Theorem 1.5. It answers a question proposed by Havet et al. in [5].
Theorem 1.5
For any positive integer $k$, there are graphs of treewidth $k$ with inversion diameter $2k$.
This paper is organized as follows. In Section 2, we give basic definitions and notation. The proofs of Theorems 1.5 and 1.4 are given in Sections 3 and 4, respectively.
2 Preliminary
Let $G=(V,E)$ be a graph. The distance between $u$ and $v$, denoted by $dist(u,v)$, is the number of edges in a shortest path joining $u$ and $v$. For any vertex $u\in V(G)$, denote $N(u)=\{v~{}|~{}uv\in E(G)\}$. Then $d(u)=|N(u)|$ is the degree of $u$. Let $\Delta=\Delta(G)$ be the maximum degree of $G$. We call $G$ $k$-regular if $d(u)=k$ for any $u\in V(G)$. Let $G$ be a graph and $S$ a vertex subset. Let $G[S]$ denote the subgraph of $G$ induced by $S$. For a graph $G$ and a vertex $v$, denote $G-v$ the graph induced by $V(G)-\{v\}$. For a graph $G$ and a induced subgraph $H$, denote $G-H$ the graph induced by $V(G)-V(H)$.
A label of $G$ is a mapping $\pi:E(G)\rightarrow\mathbb{F}_{2}$. A $t$-dim vector assignment of $G$ with the label $\pi$ is a mapping $f:V(G)\rightarrow\mathbb{F}_{2}^{t}$ such that $\pi(uv)=f(u)\cdot f(v)$ for every edge $uv\in E(G)$, where $f(u)\cdot f(v)$ to be the scalar product of $f(u)$ and $f(v)$ over $\mathbb{F}_{2}^{t}$. Usually, we use bold letter $\mathbf{u}$ to represent $f(u)$. We use $\mathbf{0}$  (resp. $\mathbf{1}$) to represent vectors in $\mathbb{F}_{2}^{t}$ whose coordinates are all 00 (resp. $1$). We say a vector $\mathbf{u}\in\mathbb{F}_{2}^{t}$ is odd (resp. even), if $\mathbf{u\cdot 1}$ is one  (resp. zero), i.e., $\mathbf{u}$ has odd (resp. even) number of $1$.
The inversion diameter has a close relationship with vector assignment as following.
Proposition 2.1 ([5])
For every graph $G$ and every positive integer $t$, the following are equivalent.
  1. 1.
    $diam(\mathcal{I}(G))\leq t$.
  2. 2.
    For every label $\pi$, there exists a $t$-dim vector assignment of $G$ with the label $\pi$.
The treewidth of a graph $G$ is denoted by $tw(G)$. There are many ways to define treewidth. Here we give a definition of treewidth from the perspective of $k$-tree.
Definition 2.2
A graph $G$ is a k-tree if
  1. 1.
    either it is a $k$-clique,
  2. 2.
    or there exists a vertex $v$ such that $N(v)$ is a $k$-clique, and $G-v$ is a k-tree.
We say a graph is a partial $k$-tree if it is a subgraph of a $k$-tree. It is known that a graph $G$ is a partial $k$-tree, if and only if the treewidth of $G$ is at most $k$ [6, 7].
Let $\mathcal{L}(\mathbf{v_{1},\ldots,v_{k}})$ denote the linear space spanned by $\mathbf{v_{1},\ldots,v_{k}}$. For two vectors $\mathbf{v}$ and $\mathbf{u}$ in $\mathbb{F}_{2}^{t}$, we write $\mathbf{v\perp u}$ if $\mathbf{v\cdot u}=0$. For a vector $\mathbf{v}\in\mathbb{F}_{2}^{t}$ and a linear space $\mathbf{U}$ in $\mathbb{F}_{2}^{t}$, we write $\mathbf{v\perp U}$ if $\mathbf{v\perp u}$ for any $\mathbf{u\in U}$. The orthogonal complementary space of $\mathbf{U}$ is $\mathbf{U}^{\perp}=\{\textbf{v}\mid\mathbf{v\perp U}\}$. For any positive integer $k$, write $[k]=\{1,2,\ldots,k\}$.
Definition 2.3
We say a sequence of vector $\mathbf{v_{1},\ldots,v_{k}}$ are orthogonal if $\mathbf{v_{i}\perp v_{j}}$ for any $i,j\in[k]$ and $i\neq j$. We say they are self-orthogonal if $\mathbf{v_{i}\perp v_{j}}$ for any $i,j\in[k]$, that is, they are orthogonal and every vector is even.
Definition 2.4
A linear space $\mathbf{U}$ is self-orthogonal if $\mathbf{U\subseteq U^{\perp}}$.
Let $\mathbf{U}$ be a self-orthogonal linear space. Then $\mathbf{U}$ is orthogonal and every vector in $\mathbf{U}$ is even. It is easy to verify that $\mathbf{U}$ is self-orthogonal if and only if it has self-orthogonal base vectors.
For a linear space $\mathbf{U}$ and a vector $\mathbf{v}$, denote $\mathbf{v+U}$ by the set $\{\mathbf{v+u}\mid\mathbf{u\in U}\}$ and denote $\mathcal{L}(\mathbf{U,v})$ the space spanned by $\mathbf{v}$ and a basis of $\mathbf{U}$, that is the summation space of $\mathbf{U}$ and $\mathcal{L}(\mathbf{v})$.
3 Proof of Theorem 1.5
For $k\geq 1$, we define a sequence of graph $G_{i}^{(k)}$ with a fixed label $\pi_{i}^{(k)}$. First, let $G_{0}^{(k)}$ be a $k$-clique with an arbitrarily label $\pi_{0}^{(k)}$. Then, we recursively construct $G_{i}^{(k)}$ as following:
(i) for each $k$-clique with vertices $v_{1},\ldots,v_{k}$ in $G_{i-1}^{(k)}$ and each $\mathbf{x}=(x_{1},\ldots,x_{k})^{T}\in\mathbb{F}_{2}^{k}$, we add a new vertex $u$ such that $uv_{j}\in E(G_{i}^{(k)})$ and $\pi_{i}^{(k)}(uv_{j})=x_{j}$, for all $1\leq j\leq k$ ;
(ii) $\pi^{(k)}_{i}|_{G^{(k)}_{i-1}}=\pi^{(k)}_{i-1}$.
Since $|\mathbb{F}_{2}^{k}|=2^{k}$, we add $2^{k}$ new vertices for each $k$-clique in $G_{i-1}^{(k)}$. By Definition 2.2, $G_{m}^{(k)}$ is a $k$-tree for any $m$, that is, of treewidth at most $k$. Since $\pi^{(k)}_{n}|_{G^{(k)}_{m}}=\pi^{(k)}_{m}$ when $n>m$, we may use $\pi^{(k)}$ to denote the label of $G^{(k)}_{m}$ for any $m$. For any vertex $v\in V(G^{(k)}_{m})$ with $m\geq 1$, there exists an unique $n$ such that $v\in V(G^{(k)}_{n})-V(G^{(k)}_{n-1})$. We say $n$ is the level of $v$, denoted by $l(v)=n$. For a vertex set $S\subseteq G^{(k)}_{m}$ with $m\geq 1$, the level of $S$ is defined to be the maximum level of vertices in $S$ denoted by $l(S)$, that is, $l(S)=\max\limits_{v\in S}\{l(v)\}$. Clearly, if $v$ is a vertex in $G_{m}^{(k)}$, then $l(v)\leq m$. Similarly, if $C$ is a vertex set in $G_{m}^{(k)}$, then $l(C)\leq m$.
Note that if $H$ is a subgraph of $G$, then $diam(\mathcal{I}(H))\leq diam(\mathcal{I}(G))$. So $(diam(\mathcal{I}(G_{m}^{(k)})))_{m\geq 0}$ is an increasing sequence with upper bound $2k$ by Theorem 1.2.
Let $\lambda^{(k)}=\lim\limits_{m\rightarrow+\infty}diam(\mathcal{I}(G_{m}^{(k)}))$. Then $\lambda^{(k)}\leq 2k$. We will show that $\lambda^{(k)}=2k$ and then $G_{m}^{(k)}$ is of inversion diameter $2k$ when $m$ is sufficiently large.
Next we suppose $\lambda^{(k)}\leq 2k-1$. Then for any $m$, $G_{m}^{(k)}$ has a $(2k-1)$-dim vector assignment with the label $\pi^{(k)}$ by Proposition 2.1. Thus for each $v\in V(G_{m}^{(k)})$, there is a vector $\mathbf{v}\in\mathbb{F}_{2}^{2k-1}$ corresponding to it. The following lemmas show the properties of the vectors assigned to $k$-cliques in $G^{(k)}_{m}$.
Lemma 3.1
If there is a $k$-clique of level $m$ with vertices $v_{1},\ldots,v_{k}$ in $G^{(k)}_{m+1}$, then $\mathbf{v_{1}},\ldots,\mathbf{v_{k}}$ are linear independent.
Proof. Otherwise, without loss of generality, assume $\mathbf{v_{1}}=\sum_{i=2}^{k}c_{i}\mathbf{v_{i}}$ where $c_{i}\in\mathbb{F}_{2}$ for all $2\leq i\leq k$. By the construction, there exists a vertex $u\in V(G^{(k)}_{m+1})$ connecting each $v_{i}$ with $\pi^{(k)}(uv_{i})=1$ for all $2\leq i\leq k$ and $\pi^{(k)}(uv_{1})=\sum_{i=2}^{k}c_{i}+1$. Therefore,
$\sum_{i=2}^{k}c_{i}+1$ $=\pi^{(k)}(uv_{1})=\mathbf{u\cdot v_{1}}=\mathbf{u}\cdot\sum_{i=2}^{k}c_{i}\mathbf{v_{i}}=\sum_{i=2}^{k}c_{i}\mathbf{u\cdot v_{i}}=\sum_{i=2}^{k}c_{i}\pi^{(k)}(uv_{i})=\sum_{i=2}^{k}c_{i},$
a contradiction. $\square$
Lemma 3.2
If there is a $k$-clique in $G^{(k)}_{m+2}$ of level $m$ with vertices $v_{1},\ldots,v_{k}$, and $u$ is a vertex of level $m+1$ connecting all ${(v_{i})}_{1\leq i\leq k}$, then either $\mathbf{v_{1},\ldots,v_{k},u}$ are linear independent, or $\mathbf{u=\sum_{i=1}^{k}v_{i}}$.
Proof. Firstly, by Lemma 3.1, $\mathbf{v_{1},\ldots,v_{k}}$ are linear independent. Note that for any $1\leq j\leq k$, $v_{1},\ldots,v_{j-1},v_{j+1},\ldots,v_{k},u$ is also a $k$-clique of level $m+1$ in $G^{(k)}_{m+2}$. Then by Lemma 3.1, for any $1\leq j\leq k$, $\mathbf{v_{1}},\ldots,\mathbf{v_{j-1}},\mathbf{v_{j+1}},\ldots,\mathbf{v_{k}},\mathbf{u}$ are linear independent.  Assume $\mathbf{u}=\sum_{i=1}^{k}c_{i}\mathbf{v_{i}}$ where $c_{i}\in\mathbb{F}_{2}$ for all $1\leq i\leq k$. If $c_{j}=0$ for some $j$, then it contradicts that $\mathbf{v_{1}},\ldots,\mathbf{v_{j-1}},\mathbf{v_{j+1}},\ldots,\mathbf{v_{k}},\mathbf{u}$ are linear independent. Therefore, $\mathbf{u=\sum_{i=1}^{k}v_{i}}$. $\square$
Lemma 3.3
Let $v_{1},\ldots,v_{k}$ be vertices of a $k$-clique of level $m$ in $G^{(k)}_{m+2}$ and $\mathbf{A}=(\mathbf{v_{1}},\ldots,\mathbf{v_{k}})^{T}$. Then for any $\mathbf{b}\in\mathbb{F}_{2}^{k}$, $\mathbf{Ax=b}$ has a solution $\mathbf{y}$ such that either $\mathbf{v_{1}},\ldots,\mathbf{v_{k}},\mathbf{y}$ are linear independent, or $\mathbf{y}=\sum_{i=1}^{k}\mathbf{v_{i}}$.
Proof. Let $\mathbf{b}=(b_{1},\ldots,b_{k})^{T}$. By the construction, there exists a vertex $u\in V(G^{(k)}_{m+1})$ of level $m+1$ connecting ${(v_{i})}_{1\leq i\leq k}$ such that $\pi^{(k)}(uv_{i})=b_{i}$ for all $1\leq i\leq k$. Then we have $\mathbf{Au=b}$. By Lemma 3.2, either $\mathbf{v_{1}},\ldots,\mathbf{v_{k}},\mathbf{u}$ are linear independent, or $\mathbf{u}=\sum_{i=1}^{k}\mathbf{v_{i}}$. $\square$
The above lemmas actually work for arbitrary $\lambda^{(k)}$. The following lemmas need the assumption $\lambda^{(k)}\leq 2k-1$. If $\mathbf{b=0}$, we have a strong conclusion.
Lemma 3.4
Let $v_{1},\ldots,v_{k}$ be vertices of a $k$-clique of level $m$ in $G^{(k)}_{m+2}$ and $\mathbf{A}=(\mathbf{v_{1}},\ldots,\mathbf{v_{k}})^{T}$. Then $\mathbf{Ax=0}$ has a solution $\mathbf{y}$ such that $\mathbf{v_{1}},\ldots,\mathbf{v_{k}},\mathbf{y}$ are linear independent.
Proof. We prove it by contradiction. By Lemma 3.1, $\mathbf{v_{1},\ldots,v_{k}}$ are linear independent. Let $\mathbf{U}$ be the solution space of $\mathbf{Ax=0}$. Suppose $\mathbf{U}$ is a subspace of $\mathcal{L}(\mathbf{v_{1}},\ldots,\mathbf{v_{k}})$. Since $\mathbf{A}$ is a $k\times(2k-1)$ matrix, $dim(\mathbf{U})=(2k-1)-k=k-1$. By letting $\mathbf{b=0}$ in Lemma 3.3, we have $\sum_{i=1}^{k}\mathbf{v_{i}}\in\mathbf{U}$.
For each $j\in[k]$, the solution set of $\mathbf{Ax=Av_{j}}$ is in $\mathbf{v_{j}+U}\subseteq\mathcal{L}(\mathbf{v_{1}},\ldots,\mathbf{v_{k}})$. By Lemma 3.3, $\sum_{i=1}^{k}\mathbf{v_{i}}\in\mathbf{v_{j}+U}$. Therefore, $\mathbf{v_{j}\in U}$ for any $1\leq j\leq k$, which contradicts that $dim(\mathbf{U})=k-1$ because $\mathbf{v_{1}},\ldots,\mathbf{v_{k}}$ are linear independent. $\square$
Definition 3.5
Let $C$ be a $p$-clique of $G^{(k)}_{m}$ for some $m$. $C$ is called a bad clique if $dim(\textbf{V}_{C}\cap\textbf{V}_{C}^{\perp})\geq p-1$, where $\textbf{V}_{C}=\mathcal{L}(\{\mathbf{v}\mid v\in C\})$ and $p\geq 1$.
Note that a single vertex is always a bad $1$-clique. If $\lambda^{(k)}\leq 2k-1$, “large” bad clique will finally cause contradictions. The following lemma is the main part of our proof which states that we can find “large” bad clique when $m$ is sufficiently large.
Lemma 3.6
If there exists a bad $p$-clique of level $m$ in $G^{(k)}_{m+k+2}$ with $p<k$, then there exists a bad clique in $G^{(k)}_{m+k+2}$ of size at least $p+1$.
Proof. We prove it by contradiction. Suppose the $p$-clique $C_{1}$ with vertices $v_{1},v_{2},\ldots,v_{p}$ of level $m$ is the largest bad clique in $G^{(k)}_{m+k+2}$, where $p<k$. Then $dim(\textbf{V}_{C_{1}})=p$ by Lemma 3.1. Let $\textbf{U}=\textbf{V}_{C_{1}}\cap\textbf{V}_{C_{1}}^{\perp}$. Then $dim(\textbf{U})\geq p-1$ by Definition 3.5. For $\phi_{1},\phi_{2}\in\textbf{U}$, we have $\phi_{1}\perp\phi_{2}$ which means U is self-orthogonal.
We first show that $dim(\textbf{U})=p-1$. Suppose $dim(\textbf{U})=p$. Then $\textbf{U}=\textbf{V}_{C_{1}}\subseteq\textbf{V}_{C_{1}}^{\perp}$. Since $p<k$, by the construction of $G^{(k)}_{m+k+2}$, there exists a vertex $u$ of level $m+1$ such that $uv_{i}\in E(G^{(k)}_{m+k+2})$ and $\pi^{(k)}(uv_{i})=0$ for each $i\in[p]$. Let $C_{2}\coloneqq{C_{1}}\cup\{u\}$. By Lemma 3.1, we have $\mathbf{u,v_{1},\ldots,v_{p}}$ are linear independent. By $\pi^{(k)}(uv_{i})=0$ for each $i\in[p]$, we have that $\mathbf{u\perp V_{C_{1}}}$. Thus $\textbf{V}_{C_{1}}\subseteq\textbf{V}_{C_{2}}\cap\textbf{V}_{C_{2}}^{\perp}$ which implies $dim(\textbf{V}_{C_{2}}\cap\textbf{V}_{C_{2}}^{\perp})\geq p$. Hence $C_{2}$ is a bad $(p+1)$-clique, a contradiction with the maximality of ${C_{1}}$. So $dim(\textbf{U})=p-1$. In fact, we conclude that U is a self-orthogonal $(p-1)$-dim subspace of $\mathbb{F}_{2}^{2k-1}$ and then each vector in U is even.
Since $dim(\textbf{U})=p-1$, there is $\textbf{v}_{i}\in\{\textbf{v}_{1},\textbf{v}_{2},\dots,\textbf{v}_{p}\}$, say $i=1$, such that $\textbf{v}_{1}\notin\textbf{U}$. Then $\mathcal{L}(\textbf{U},\textbf{v}_{1})=\textbf{V}_{C_{1}}$. If $\mathbf{v_{1}}$ is even, then $\mathbf{v_{1}}\perp\mathcal{L}(\textbf{U},\textbf{v}_{1})$, which contradicts with $\mathbf{v_{1}}\notin\textbf{U}$. Thus we have $\textbf{v}_{1}$ is odd.
Claim 3.7
If $u$ is a vertex in $G^{(k)}_{m+k}$ such that $uv_{i}\in E(G^{(k)}_{m+k})$ and $\pi^{(k)}(uv_{i})=0$ for each $i\in[p]$, then u is odd.
[…]
Fix a vertex $v_{0}$ of level $m+1$ such that $v_{0}v_{i}\in E(G^{(k)}_{m+k})$ and $\pi^{(k)}(v_{0}v_{i})=0$ for each $i\in[p]$. Then $\mathbf{v}_{0}\perp\mathbf{V_{C_{1}}}$. By the construction of $G^{(k)}_{m+k}$, there exists a set of vertices $C_{2}=\{w_{1},w_{2},\dots,w_{k-p}\}\subseteq V(G^{(k)}_{m+k})$ satisfying the following:
  1. 1.
    $\{v_{0},v_{1},\ldots,v_{p},w_{1},w_{2},\ldots,w_{k-p}\}$ is a $(k+1)$-clique;
  2. 2.
    $\pi^{(k)}(w_{i}w_{j})=0$, for each $i,j\in[k-p],i\neq j$;
  3. 3.
    and $\pi^{(k)}(v_{i}w_{j})=\textbf{v}_{i}\cdot(\textbf{v}_{1}+\textbf{v}_{0})$, for each $i\in[0,p],j\in[k-p]$.
Claim 3.8
$\textbf{w}_{j}$ is even for each $j\in[k-p]$.
Proof of Claim 3.8. Suppose there is $j\in[k-p]$ such that $\textbf{w}_{j}$ is odd. Let $\beta\coloneqq\textbf{w}_{j}+\textbf{v}_{1}$. Recall that $\textbf{v}_{1}$ is odd. Then
$\beta\cdot\textbf{w}_{j}=\textbf{w}_{j}\cdot\textbf{w}_{j}+\textbf{v}_{1}\cdot\textbf{w}_{j}=\textbf{w}_{j}\cdot\textbf{w}_{j}+\textbf{v}_{1}\cdot(\textbf{v}_{1}+\textbf{v}_{0})=0.$
On the other hand, for any $v_{i}\in C_{1}$,
$\beta\cdot\textbf{v}_{i}=\textbf{w}_{j}\cdot\textbf{v}_{i}+\textbf{v}_{1}\cdot\textbf{v}_{i}=\textbf{v}_{i}\cdot\textbf{v}_{1}+\textbf{v}_{i}\cdot\textbf{v}_{0}+\textbf{v}_{1}\cdot\textbf{v}_{i}=0.$
Hence $\beta\perp\textbf{V}_{C_{1}\cup\{w_{j}\}}$. We have $\beta\in\textbf{V}_{C_{1}\cup\{w_{j}\}}\cap\textbf{V}_{C_{1}\cup\{w_{j}\}}^{\perp}$ and $\textbf{U}\subseteq\textbf{V}_{C_{1}\cup\{w_{j}\}}\cap\textbf{V}_{C_{1}\cup\{w_{j}\}}^{\perp}$. From Lemma 3.1 and $dim(\textbf{U})=p-1$, we know $\textbf{w}_{j}\notin\textbf{U}$. Then $dim(\textbf{V}_{C_{1}\cup\{w_{j}\}}\cap\textbf{V}_{C_{1}\cup\{w_{k}\}}^{\perp})\geq p$ which implies $C_{1}\cup\{w_{j}\}$ is a bad $(p+1)$-clique, a contradiction with the maximality of $C_{1}$.
$\textbf{v}_{0}$ $\textbf{v}_{1}$ $\textbf{v}_{i}$ $\textbf{w}_{j_{1}}$ $\textbf{w}_{j_{2}}$ $\beta_{j_{1}}$ $\beta_{j_{2}}$
$\textbf{v}_{0}$ 1 0 0 1 1 0 0
$\textbf{v}_{1}$ 0 1 $\textbf{v}_{1}\cdot\textbf{v}_{i}$ 1 1 0 0
$\textbf{v}_{i}$ 0 $\textbf{v}_{1}\cdot\textbf{v}_{i}$ $\textbf{v}_{i}\cdot\textbf{v}_{i}$ $\textbf{v}_{1}\cdot\textbf{v}_{i}$ $\textbf{v}_{1}\cdot\textbf{v}_{i}$ 0 0
$\textbf{w}_{j_{1}}$ 1 1 $\textbf{v}_{1}\cdot\textbf{v}_{i}$ 0 0 0 0
$\textbf{w}_{j_{2}}$ 1 1 $\textbf{v}_{1}\cdot\textbf{v}_{i}$ 0 0 0 0
$\beta_{j_{1}}$ 0 0 0 0 0 0 0
$\beta_{j_{2}}$ 0 0 0 0 0 0 0
Table 1: The inner products for $i\in[p]$ and $j_{1},j_{2}\in[k-p]$
Now we complete the proof of Lemma 3.6. For each $j\in[k-p]$, let $\beta_{j}=\textbf{v}_{0}+\textbf{v}_{1}+\textbf{w}_{j}$. By Claim 3.7, $\textbf{v}_{0}$ is odd. By Claim 3.8, $\textbf{w}_{j}$ is even for each $j\in[k-p]$. It is not difficult to show that $\beta_{j}\perp\textbf{v}_{0}$, $\beta_{j}\perp\textbf{V}_{C_{1}}$ and $\beta_{j}\perp\textbf{V}_{C_{2}}$. Check Table 1 for the inner products between the vectors that we are working on. Let $C_{3}=C_{1}\cup C_{2}$ and $W=\textbf{V}_{C_{3}}=\textbf{V}_{C_{1}}+\textbf{V}_{C_{2}}$. Then $dim(W)=k$ from Lemma 3.1. Since $W\subseteq\mathbb{F}_{2}^{2k-1}$, we have $dim(W^{\perp})=k-1$. Let $W^{\prime}=\mathcal{L}(\textbf{U},\beta_{1},\dots,\beta_{k-p})$. Since $\textbf{U}\perp W$ and $\beta_{j}\perp W$ for each $j\in[k-p]$, we have $W^{\prime}\subseteq W^{\perp}$. Note that $\textbf{V}_{C_{1}},\textbf{V}_{C_{2}}\subseteq\mathcal{L}(\textbf{v}_{0},\textbf{v}_{1},W^{\prime})$. We have $W\subseteq\mathcal{L}(\textbf{v}_{0},\textbf{v}_{1},W^{\prime})$ which implies $dim(W^{\prime})\geq k-2$. If $\textbf{v}_{0}\notin W$, then $dim(W^{\prime})\geq k-1$ which implies $W^{\perp}=W^{\prime}$. Since $\textbf{v}_{0}\perp W^{\prime}$, we have $\textbf{v}_{0}\perp W^{\perp}$, a contradiction with $\textbf{v}_{0}\notin W$. Hence $\textbf{v}_{0}\in W$. By Lemma 3.2, we have $\textbf{v}_{0}+\dots+\textbf{v}_{p}+\textbf{w}_{1}+\dots+\textbf{w}_{k-p}=0$.
Since $\textbf{v}_{0}\in W$, we have $\beta_{j}\in W$ for each $j$ which implies $W^{\prime}\subseteq W$. So $W^{\prime}\subseteq W\cap W^{\perp}$. If $dim(W^{\prime})\geq k-1$, then $C_{3}$ is a bad $k$-clique, a contradiction with the maximality of $C_{1}$. Hence $dim(W^{\prime})=k-2$. Let $\alpha\in W^{\perp}\backslash W^{\prime}$ such that $W^{\perp}=\mathcal{L}(W^{\prime},\alpha)$. By the construction of $G^{(k)}_{m+k}$, there exists a vertex $x$ connecting to all vertices of $C_{3}$ such that $\pi^{(k)}(xy)=0$ for each $y\in C_{3}$. Then $\textbf{x}\in W^{\perp}$. From Claim 3.7, x is odd. Since U is a self-orthogonal subspace and $\beta_{i}\bot\beta_{j}$ for any $i,j\in[k-p]$ (see Table 1), we have that the vectors in $W^{\prime}$ are all even. By $\textbf{x}\in W^{\perp}=\mathcal{L}(W^{\prime},\alpha)$ and x being odd, we have $\alpha$ is odd.
Let $C^{*}=\{v_{0},\ldots,v_{p},w_{1},\ldots,w_{k-p-1}\}$ (if $p=k-1$, then let $C^{*}=\{v_{0},\ldots,v_{p}\}$). Then there is $x^{*}$ connecting to all vertices of $C^{*}$ such that $\pi^{(k)}(x^{*}y)=\textbf{v}_{0}\cdot\textbf{y}$ for each $y\in C^{*}$. Since $\textbf{v}_{0}+\dots+\textbf{v}_{p}+\textbf{w}_{1}+\dots+\textbf{w}_{k-p}=0$, by Lemma 3.1, $W=\textbf{V}_{C^{*}}$. Then $\textbf{x}^{*}\in\textbf{v}_{0}+W^{\perp}$. Since $\pi^{(k)}(x^{*}v_{i})=\textbf{v}_{0}\cdot\textbf{v}_{i}=0$ for each $i\in[k]$, from Claim 3.7, $\textbf{x}^{*}$ is odd. Note that $\textbf{x}^{*}\in\textbf{v}_{0}+\mathcal{L}(\alpha,W^{\prime})$. Since all vectors in $W^{\prime}$ are even and $\textbf{v}_{0},\alpha$ are odd, we have $\textbf{x}^{*}\in\textbf{v}_{0}+W^{\prime}\subseteq W$. From Lemma 3.2, $\textbf{x}^{*}=\textbf{v}_{0}+\dots+\textbf{v}_{p}+\textbf{w}_{1}+\dots+\textbf{w}_{k-p-1}=\textbf{w}_{k-p}$. Since $\textbf{x}^{*}\cdot\mathbf{v_{1}}=0\neq 1=\textbf{w}_{k-p}\cdot\mathbf{v_{1}}$ , we derive a contradiction. $\square$
With the help of those lemmas, we can derive a contradiction when $\lambda^{(k)}\leq 2k-1$ and hence, $\lambda^{(k)}=2k$.
Theorem 3.9
$\lambda^{(k)}=2k$.
Proof. Suppose $\lambda^{(k)}\leq 2k-1$. By Lemma 3.6, the largest bad clique $C_{0}$ in $G^{(k)}_{k(k+2)}$ is of size $k$. Then $dim(\textbf{V}_{C_{0}}^{\perp}\cap\textbf{V}_{C_{0}})\geq k-1$. By Lemma 3.1, $dim(\textbf{V}_{C_{0}})=k$ and then $dim(\textbf{V}_{C_{0}}^{\perp})=k-1$ by $\textbf{V}_{C_{0}}\subseteq\mathbb{F}_{2}^{2k-1}$. We have $\textbf{V}_{C_{0}}^{\perp}\subseteq\textbf{V}_{C_{0}}$ by checking the dimensions. Then we can derive a contradiction by Lemma 3.4. $\square$
Now we can give the proof of our main Theorem.
Proof of Theorem 1.5. For any $k\geq 1$, we have $\lambda^{(k)}=2k$ by Theorem 3.9. Then there exists a $M_{k}$ such that for every $m\geq M_{k}$, $diam(\mathcal{I}(G^{(k)}_{m}))=2k$. Thus for all $m\geq M_{k}$, $G^{(k)}_{m}$ are desired graphs of treewidth at most $k$ and inversion diameter $2k$. $\square$
Figure 1: An example of outer-planar graph with labeled edges of inversion diameter $4$ verified by computer.
Note that any outer-planar graph is of treewidth $2$ and hence has inversion diameter at most $4$ by Lemma 1.2. We construct an outer-planar graph of inversion diameter $4$ verified by computer as Figure 1. The codes are available on GitHub. Therefore, the upper bound $diam(\mathcal{I}(G))\leq 4$ for any outer-planar graph $G$ is tight.
4 Proof of Theorem 1.4
In this section, we intend to give the proof of Theorem 1.4.
Let $G$ be a graph. We say $G$ is $3$-critical if $diam(\mathcal{I}(G))>3$ and for any proper subgraph $G^{\prime}$, $diam(\mathcal{I}(G^{\prime}))\leq 3$. Clearly, a $3$-critical graph is connected. If $G$ is $3$-critical, by Proposition 2.1, there exists a label $\pi$ such that there is no $3$-dim vector assignment of $G$ with $\pi$. We call such $\pi$ a bad label.
Let $G$ be a $3$-critical graph with a bad label $\pi$ and $H$ an induced subgraph of $G$. Denote $N_{G}(H)=\{v\in V(G)-V(H)\mid\exists u\in V(H),uv\in E(G)\}$. By the definition of $3$-critical graph, $G-H$ admits a $3$-dim vector assignment $f:V(G-H)\rightarrow\mathbb{F}_{2}^{3}$ with $\pi|_{G-H}$. For a vertex $v\in N_{G}(H)$, define $\mathcal{A}_{f}(v)=\{\mathbf{v}\in\mathbb{F}_{2}^{3}\mid\mathbf{v\cdot}f(u)=\pi(uv),\mbox{~{}for any~{}}uv\in E(G-H)\}$. Note that $f(v)\in\mathcal{A}_{f}(v)$.
Let $H$ be a subgraph of $G$. We say $(\mathcal{B}_{f}(v))_{v\in N_{G}(H)}$ is an available boundary family if we can assign each vertex $v\in N_{G}(H)$ a set $\mathcal{B}_{f}(v)$ such that:
  1. 1.
    $f(v)\in\mathcal{B}_{f}(v)\subseteq\mathcal{A}_{f}(v)$, and
  2. 2.
    $\{v\in N_{G}(H)\mid|\mathcal{B}_{f}(v)|\geq 2\}$ is an independent set in $G-H$.
When there is no ambiguity, we may ignore the subscript $f$.
The following lemma states that if we already have a vector assignment of $G-H$, then we can reassign the vectors for $v\in N_{G}(H)$ from $\mathcal{B}(v)$ and the result is also a valid vector assignment.
Lemma 4.1
Let $H$ be an induced subgraph of a $3$-critical graph $G$ with a bad label $\pi$. Let $f$ be a $3$-dim vector assignment on $G-H$ with $\pi|_{G-H}$ and $(\mathcal{B}_{f}(v))_{v\in N_{G}(H)}$ an available boundary family. Then for any $3$-dim vector assignment $g$ on $G-H$ satisfying
  1. 1.
    $g(v)=f(v)$, $\forall v\in V(G-H)-N_{G}(H)$, and
  2. 2.
    $g(v)\in\mathcal{B}_{f}(v)$, $\forall v\in N_{G}(H)$,
we have $g$ is a $3$-dim vector assignment on $G-H$ with $\pi|_{G-H}$.
Proof. We only need to verify that $g(v)\cdot g(u)=\pi(uv)$ for all $uv\in E(G-H)$. Note that $A\coloneqq\{v\in V(G-H)\mid g(v)\neq f(v)\}\subseteq\{v\in N_{G}(H)\mid|\mathcal{B}_{f}(v)|\geq 2\}$ from the definition. Then $\{v\in V(G-H)\mid g(v)\neq f(v)\}$ is an independent set. Since we already have $f(v)\cdot f(u)=\pi(uv)$ for all $uv\in E(G-H)$ and $\{v\in V(G-H)\mid g(v)\neq f(v)\}$ is an independent set, we now only need to verify that $g(v)\cdot g(u)=\pi(uv)$ for all $uv\in E(G-H)$ satisfying $u\in A$ and $v\notin A$. Since $g(u)\in\mathcal{B}_{f}(u)$ and $g(v)=f(v)$, we have $g(v)\cdot g(u)=\pi(uv)$ by the definition of $\mathcal{B}_{f}(u)$. $\square$
We say $H$ is reducible if there exists an available boundary family $(\mathcal{B}_{f}(v))_{v\in N_{G}(H)}$ and a $3$-dim vector assignment $g$ on $G[V(H)\cup N_{G}(H)]$ with $\pi|_{G[V(H)\cup N_{G}(H)]}$ such that $g(v)\in\mathcal{B}_{f}(v)$ for any $v\in N_{G}(H)$. The following lemma states that there is no reducible structure in $3$-critical graph.
Lemma 4.2
Let $G$ be a $3$-critical graph with a bad label $\pi$. Then there is no reducible induced subgraph of $G$.
Proof. Suppose $H$ is an induced reducible subgraph of $G$. Then $G-H$ admits a $3$-dim vector assignments $f$ with $\pi|_{G-H}$, an available boundary family $(\mathcal{B}_{f}(v))_{v\in N_{G}(H)}$ and a $3$-dim vector assignment $g$ on $G[V(H)\cup N_{G}(H)]$ such that $g(v)\in\mathcal{B}_{f}(v)$ for any $v\in N_{G}(H)$. Define a function $h:V(G)\rightarrow\mathbb{F}_{2}^{3}$ by letting $h(v)=f(v)$ for any $v\in V(G-H)-N_{G}(H)$ and $h(v)=g(v)$ for any $v\in N_{G}(H)\cup V(H)$. By the definition, $h|_{G[V(H)\cup N_{G}(H)]}$ is a $3$-dim vector assignment with label $\pi|_{G[V(H)\cup N_{G}(H)]}$. By Lemma 4.1, $h|_{G-H}$ is a $3$-dim vector assignment of $G-H$ with label $\pi|_{G-H}$. Since there is no edge between $V(G-H)-N_{G}(H)$ and $V(H)$, $h$ is a $3$-dim vector assignment of $G$ with $\pi$, a contradiction. $\square$
In the following, we are going to find certain reducible structures in $3$-critical graph.
Lemma 4.3
Let $G$ be a $3$-critical graph with a bad label $\pi$. For any vertex $v\in V(G)$, at least one edge adjacent to $v$ is labeled $1$ by $\pi$.
Proof. Suppose there exists a vertex $v\in V(G)$ such that $\pi(uv)=0$ for all $u\in N_{G}(v)$. Let $G^{\prime}=G-v$. Then $G^{\prime}$ admits a $3$-dim vector assignment $f$ with $\pi|_{G^{\prime}}$. Let $f(v)=\mathbf{0}\in\mathbb{F}_{2}^{3}$. Then it is not difficult to verify that $f$ is a $3$-dim vector assignment of $G$ with $\pi$, a contradiction. $\square$
Lemma 4.4
Let $G$ be a graph with a label $\pi$. If $G$ admits a $3$-dim vector assignment with $\pi$, then there exists a $3$-dim vector assignment $f$ with $\pi$ such that for every vertex $v\in V(G)$ of degree at most $2$, $f(v)\neq\mathbf{0}$.
Proof. Let $f$ be the $3$-dim vector assignment of $G$ with $\pi$ which minimizes $n_{f}=|\{v\in V(G)\mid f(v)=\mathbf{0},d_{G}(v)\leq 2\}|$. Suppose $n_{f}>0$. Let $w\in\{v\in V(G)\mid f(v)=\mathbf{0},d_{G}(v)\leq 2\}$ and $\mathcal{A}(w)=\{\mathbf{w}\in\mathbb{F}_{2}^{3}\mid\mathbf{w\cdot}f(u)=\pi(uw),\forall uw\in E(G)\}$. Then $|\mathcal{A}(w)|\geq 2$ since $d_{G}(w)\leq 2$. Choose $\mathbf{w}\in\mathcal{A}(w)-\{\mathbf{0}\}$ and define a function $g:V(G)\rightarrow\mathbb{F}_{2}^{3}$ by letting $g(v)=f(v)$ for any $v\in V(G)-\{w\}$ and $g(w)=\mathbf{w}$. It is easy to verify that $g$ is a $3$-dim vector assignment of $G$ with $\pi$, but $n_{g}<n_{f}$, a contradiction. $\square$
Lemma 4.5
Let $G$ be a $3$-critical graph of maximum degree $3$ with a bad label $\pi$. Then $G$ is 3-regular.
Proof. Suppose there exists a vertex $v\in V(G)$ such that $d(v)=1$. Let $uv\in E(G)$. Then by Lemma 4.3, $\pi(uv)=1$. Let $V(H)=\{v\}$. Then $N_{G}(H)=\{u\}$. By hypothesis, $G-v$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-v}$. Since $d_{G-v}(u)\leq 2$, $|\mathcal{A}_{f}(u)|\geq 2$. Let $\mathcal{B}_{f}(u)=\mathcal{A}_{f}(u)$. Then $(\mathcal{B}_{f}(u))_{u\in N_{G}(H)}$ is an available boundary family. Let $g(u)\in\mathcal{B}(u)-\{\mathbf{0}\}$ and we can choose $g(v)\in\mathbb{F}_{2}^{3}$ such that $g(v)\cdot g(u)=1$. Then $H$ is reducible, a contradiction with Lemma 4.2.
Suppose there exists a vertex $v\in V(G)$ such that $d(v)=2$. Let $N_{H}(v)=\{u_{1},u_{2}\}$. By Lemma 4.3, without loss of generality, assume $\pi(vu_{1})=1$. Let $V(H)=\{v\}$. Then $N_{G}(H)=\{u_{1},u_{2}\}$. By hypothesis and Lemma 4.4, $G-v$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-v}$ such that $f(u_{1}),f(u_{2})\neq\mathbf{0}$. Let $\mathcal{B}(u_{1})=\{f(u_{1})\}$ and $\mathcal{B}(u_{2})=\mathcal{A}(u_{2})$. Then $(\mathcal{B}(u_{i}))_{i=1,2}$ is an available boundary family. Since $d_{G-v}(u_{2})\leq 2$, we have $|\mathcal{B}(u_{2})|\geq 2$. Let $g(u_{1})=f(u_{1})$. If $\pi(vu_{2})=1$ (resp. $\pi(vu_{2})=0$), choose $g(u_{2})\in\mathcal{B}(u_{2})-\{\mathbf{0}\}$ (resp. $g(u_{2})\in\mathcal{B}(u_{2})-\{f(u_{1})\}$). It is easy to verify in either case, there exists $g(v)\in\mathbb{F}_{2}^{3}$ such that $g(v)\cdot g(u_{i})=\pi(vu_{i})$ for $i=1,2$. So $H$ is reducible, a contradiction with Lemma 4.2. $\square$
Figure 2: $K_{4}^{-}$ in $G$.
Figure 3: Triangle in $G$.
Lemma 4.6
Let $G$ be a $3$-critical 3-regular graph with a bad label $\pi$. There is no induced $K_{4}^{-}$ in $G$, where $K_{4}^{-}$ is the graph obtained by deleting an edge in $K_{4}$.
Proof. Suppose there exists a $K_{4}^{-}$ in $G$ with vertex set $\{v_{0},v_{1},u_{0},u_{1}\}$ and $u_{0}u_{1}\notin E(G)$ (See Figure 3).
Let $H=G[\{v_{0},v_{1}\}]$. Then $N_{G}(H)=\{u_{0},u_{1}\}$. By hypothesis and Lemma 4.4, $G-H$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-H}$ such that $f(u_{0}),f(u_{1})\neq\mathbf{0}$. Let $\mathcal{B}(u_{i})=\mathcal{A}_{f}(u_{i}),i=0,1$. Then $(\mathcal{B}(u_{i}))_{i=0,1}$ is an available boundary family. We have the following properties:
  1. 1.
    For each $i\in\{0,1\}$, $|\mathcal{B}(u_{i})|\geq 4$ by $d_{G-H}(u_{i})=1$.
  2. 2.
    For each $i\in\{0,1\}$, if $\pi(v_{0}u_{i})=\pi(v_{1}u_{i})=0$, then $\mathbf{0}\notin\mathcal{B}(u_{i})$ by Lemma 4.3.
  3. 3.
    For each $i\in\{0,1\}$, at least one edge in $\{v_{0}v_{1},v_{i}u_{0},v_{i}u_{1}\}$ is labeled one by Lemma 4.3.
With above properties, we claim that $H$ is reducible. The claim is proved by computer by enumerating all available boundary families with above properties. The source codes can be found on GitHub. Therefore, we derive a contradiction with Lemma 4.2. $\square$
Lemma 4.7
Let $G$ be a $3$-critical 3-regular graph with a bad label $\pi$. Then there is no triangle in $G$.
Proof. Suppose there exists a triangle with vertices $\{v_{0},v_{1},v_{2}\}$ and let $u_{i}$ be the neighbor of $v_{i}$, for $i=0,1,2$ (See Figure 3). By Lemma 4.6, $u_{0},u_{1},u_{2}$ are either distinct vertices, or $u_{0}=u_{1}=u_{2}$. If $u_{0}=u_{1}=u_{2}$, then $G=K_{4}$ by $G$ being 3-regular. However, $diam(\mathcal{I}(K_{4}))=3$ [5], which contradicts that $G$ is $3$-critical. Hence, we conclude that $u_{0},u_{1},u_{2}$ are distinct vertices. Let $V(H)=\{v_{0},v_{1},v_{2}\}$. Then $N_{G}(H)=\{u_{0},u_{1},u_{2}\}$. By hypothesis and Lemma 4.4, $G-H$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-H}$ such that $f(u_{i})\neq\mathbf{0},i=0,1,2$. We can assume, without loss of generality, that $u_{0}$ satisfies the property: if $f(u_{1})=f(u_{2})$, then $f(u_{0})=f(u_{1})=f(u_{2})$. Let $\mathcal{B}(u_{0})=\mathcal{A}(u_{0})$ and $\mathcal{B}(u_{i})=\{f(u_{i})\},i=1,2$. Now we have the following properties:
  1. 1.
    $|\mathcal{B}(u_{0})|\geq 2$ by $d_{G-H}(u_{0})=2$.
  2. 2.
    For each $i=0,1,2$, at least one edge adjacent to $v_{i}$ is labeled one by Lemma 4.3.
  3. 3.
    If $\pi(u_{0}v_{0})=0$, then $\mathbf{0}\notin\mathcal{B}(u_{0})$, also by Lemma 4.3.
  4. 4.
    If $f(u_{1})=f(u_{2})$, then $f(u_{1})=f(u_{0})\in\mathcal{B}(u_{0})$.
With above properties, we claim that $H$ is reducible which is proved by computer. The source codes can be found on GitHub. Therefore, we derive a contradiction with Lemma 4.2. $\square$
Figure 4: $P_{3}$ with edges labeled one in $G$.
Figure 5: $K_{2,3}$ in $G$.
Lemma 4.8
Let $G$ be a $3$-critical 3-regular graph with a bad label $\pi$. Then there is no $P_{3}$ with two edges labeled one in $G$.
Proof. Suppose there exists a path $wu_{0}u^{\prime}$ such that $\pi(wu_{0})=\pi(u_{0}u^{\prime})=1$. By Lemma 4.7, $u^{\prime}w\notin E(G)$. Let $u_{1},u_{2}$ be the neighbor of $w$ (See Figure 5). By Lemma 4.7, $\{u_{0},u_{1},u_{2}\}$ is an independent set. Let $V(H)=\{w\}$. Then $N_{G}(H)=\{u_{0},u_{1},u_{2}\}$. By hypothesis, $G-H$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-H}$. Let $\mathcal{B}(u_{i})=\mathcal{A}(u_{i}),i=0,1,2$. Then $(\mathcal{B}(u_{i}))_{i=0,1,2}$ is an available boundary family. We have the following properties:
  1. 1.
    For each $i\in\{0,1,2\}$, $|\mathcal{B}(u_{i})|\geq 2$ by $d_{G-H}(u_{i})=2$.
  2. 2.
    $\mathbf{0}\notin\mathcal{B}(u_{0})$, because $\pi(u_{0}u^{\prime})=1$.
  3. 3.
    For each $i=1,2$, if $\pi(wu_{i})=0$, then $\mathbf{0}\notin\mathcal{B}(u_{i})$ by Lemma 4.3.
With above properties, we claim that $H$ is reducible which is proved by computer (GitHub). Therefore, we derive a contradiction with Lemma 4.2. $\square$
Lemma 4.9
Let $G$ be a $3$-critical 3-regular graph with a bad label $\pi$. Then there is no $K_{2,3}$ in $G$.
Proof. Suppose there exists a $K_{2,3}$ with vertices $\{v_{i}\}_{i=0,1}\cup\{u_{i}\}_{i=0,1,2}$ and $u_{i}v_{j}\in E(G)$ for every $i=0,1$ and $j=0,1,2$ (See Figure 5). By Lemmas 4.3 and 4.8, without loss of generality, we can assume $\pi(v_{0}u_{0})=\pi(v_{1}u_{2})=1$ and other edges in $K_{2,3}$ are labeled zero. By Lemma 4.7, $\{u_{i}\}_{i=0,1,2}$ is an independent set. Let $H=G[\{v_{0},v_{1}\}]$. Then $N_{G}(H)=\{u_{0},u_{1},u_{2}\}$. By hypothesis, $G-H$ admits a $3$-dim vector assignment $f$. Let $\mathcal{B}(u_{i})=\mathcal{A}(u_{i}),i=0,1,2$. Then $(\mathcal{B}(u_{i}))_{i=0,1,2}$ is an available boundary family. We have the following properties:
  1. 1.
    For each $i\in\{0,1,2\}$, $|\mathcal{B}(u_{i})|\geq 4$ by $d_{G-H}(u_{i})=1$.
  2. 2.
    By Lemma 4.3, $\mathbf{0}\notin\mathcal{B}(u_{1})$ and $\mathbf{0}\in\mathcal{B}(u_{i})$ for $i=0,2$.
With above properties, we claim that $H$ is reducible which is proved by computer (GitHub). Therefore, we derive a contradiction with Lemma 4.2. $\square$
Figure 6: $C_{4}$ with at least one edge labeled one in $G$.
Figure 7: Edge labeled one in $G$.
Lemma 4.10
Let $G$ be a $3$-critical 3-regular graph with a bad label $\pi$. Then there is no $C_{4}$ with at least on edge is labeled one in $G$.
Proof. Suppose there exists a $C_{4}$ with vertices $\{v_{i}\}_{i=0,1,2,3}$ and $\pi(v_{0}v_{1})=1$. Let $u_{i}$ be the neighbor of $v_{i}$ for $i=0,1,2,3$ (See Figure 7). By Lemmas 4.7 and 4.9, $\{u_{i}\}_{i=0,1,2,3}$ are distinct vertices. By Lemma 4.8, $\pi(v_{0}v_{2})=\pi(v_{1}v_{3})=0$. Let $H=C_{4}$. Then $N_{G}(H)=\{u_{0},u_{1},u_{2},u_{3}\}$. By hypothesis and Lemma 4.4, $G-H$ admits a $3$-dim vector assignment $f$ with $\pi|_{G-H}$ such that $f(u_{i})\neq\mathbf{0},i=0,1,2,3$.
Let us first consider the case $\pi(v_{2}v_{3})=0$. In this case, let $\mathcal{B}(u_{i})=\{f(u_{i})\},i=0,1,2,3$. Then $(\mathcal{B}(u_{i}))_{i=0,1,2,3}$ is an available boundary family. We claim $H$ is reducible with $(\mathcal{B}(u_{i}))_{i=0,1,2,3}$ which is proved by computer (GitHub), a contradiction with Lemma 4.2.
Now suppose $\pi(v_{2}v_{3})=1$. Then $\pi(v_{i}u_{i})=0,i=0,1,2,3$ by Lemma 4.8. Since $G$ is 3-regular, there exists $t\in\{1,2,3\}$ such that $u_{0}u_{t}\notin E(G)$. Let $\mathcal{B}(u_{i})=\mathcal{A}(u_{i})$ for $i=0,t$ and $\mathcal{B}(u_{i})=\{f(u_{i})\}$ for other $i$. Then $(\mathcal{B}(u_{i}))_{i=0,1,2,3}$ is an available boundary family. Since $\pi(v_{i}u_{i})=0$ for each $i$, we have $\mathbf{0}\notin\mathcal{B}(u_{i})$ by Lemma 4.3. Moreover, for $i=0,t$, we have $|\mathcal{B}(u_{i})|\geq 2$ since $d_{G-H}(u_{i})=2$. We claim that we can choose $\mathbf{u_{i}}\in\mathcal{B}(u_{i})$ for any $i\in\{0,1,2,3\}$ such that $\mathbf{u_{j_{0}}=u_{j_{1}}}$ and $\mathbf{u_{j_{2}}=u_{j_{3}}}$ do not occur, where $\{j_{0},j_{1},j_{2},j_{3}\}=\{0,1,2,3\}$. The claim can be proved by simple discussion. Then define $g:V(G-H)\rightarrow\mathbb{F}_{2}^{3}$ by letting $g(u_{i})=\mathbf{u_{i}},i=0,1,2,3$ and $g(v)=f(v)$ for all other vertex $v$. By lemma 4.1, $g$ is a $3$-dim vector assignment of $G-H$ with $\pi|_{G-H}$.
Let $\mathcal{B}_{g}(u_{i})=g(u_{i}),i=0,1,2,3$, then $(\mathcal{B}_{g}(u_{i}))_{i=0,1,2,3}$ is an available boundary family. We claim $H$ is reducible with $(\mathcal{B}_{g}(u_{i}))_{i=0,1,2,3}$ which is proved by computer (GitHub). A contradiction. $\square$
Now we have plenty of forbidden structures in $G$, and we can finally prove Theorem 1.4.
Proof of Theorem 1.4. By contradiction. Let $G$ be the counter example with minimum number of vertices and then minimum edges. Then $G$ is $3$-critical. Since $\Delta(G)\leq 3$, $G$ is 3-regular by Lemma 4.5. Let $\pi$ be a bad label of $G$.
We assume $v_{0}v_{1}\in E(G)$ satisfying $\pi(v_{0}v_{1})=1$ by Lemma 4.3. Let $u_{0},u_{1}$ be the neighbor of $v_{0}$ and $u_{2},u_{3}$ the neighbor of $v_{1}$ (See Figure 7). By Lemmas 4.7 and 4.10, $\{u_{i}\}_{1,2,3,4}$ is an independent set. Let $H=G[\{v_{0},v_{1}\}]$. Then $N_{G}(H)=\{u_{0},u_{1},u_{2},u_{3}\}$. By hypothesis, $G-H$ admits a $3$-dim vector assignment $f$. Let $\mathcal{B}(u_{i})=\mathcal{A}(u_{i}),i=0,1,2,3$. Then $(\mathcal{B}(u_{i}))_{i=0,1,2,3}$ is an available boundary family. Since $d_{G-H}(u_{i})=2,i=0,1,2,3$, we have $|\mathcal{B}(u_{i})|\geq 2,i=0,1,2,3$. By Lemma 4.4, we have $\mathbf{0}\notin\mathcal{B}(u_{i})$ for every $i=0,1,2,3$. Then we claim that $H$ is reducible which is proved by computer (GitHub), a contradiction with Lemma 4.2. $\square$
Acknowledgement
Y. Yang is supported by the Fundamental Research Funds for the Central University (Grant 500423306) in China. M. Lu is supported by the National Natural Science Foundation of China (Grant 12171272 & 12161141003).